FEL Upgrade Project Weekly Brief
January 27-31, 2003

We complete an extremely busy month of January with FEL Upgrade installation and commissioning activities. This week: we had the gun up to 240 kV during its HV conditioning exercise, the first recirculation arc installation was completed except for one more dipole magnet (which is scheduled for next week); all the instrumentation and controls for the high reflector optical cavity were successfully exercised.

Commissioning Activities:

1) After the second bakeout was completed last weekend, the gun vacuum improved to the 1x10e-11 torr range. High voltage conditioning took the gun to 240 kV when the process was interrupted due to a leak which developed on or around the large conflat flange which holds the ceramic assembly. The leak area was localized on Friday and the gun system will be examined from the inside on Monday after the clean hoods have cleansed the gun area over the weekend.

2) The second GQ and first GX magnet and vacuum chamber were installed in the first recirculation arc this week. (See photo below) Thus only one GX magnet assembly remains to be installed before the first recirculation arc is complete. This assembly is scheduled for installation next week.

3) The first light diagnostic magnet and dump assembly including local shielding material was completed this week. (See photo below).

4) All the instrumentation and controls for the high reflector optical cavity were successfully checked out. The final hot nitrogen purge for this assembly prior to going under vacuum was being prepared by week’s end.

5) 90% of the hardware for the Machine Protection System has been checked out.

6) Two of the five trim rack power supplies for the magnet power systems have passed their qualification tests.

  Installation in the first recirculation arc                                      1st light diagnostic magnet & dump assembly

Chris Behre (L) and Joe Gubeli (R) pose with the outcoupler assembly just before it's installation in the vacuum vessel.
The first light cavity optic is the orange object in the assembly.

G. Neil was invited to make a presentation to the DoD Joint Technology Office on Feb. 25th on the proposal we submitted to the JTO in December on investigating high power optical damage effects with the FEL Upgrade.

We prepared a status report on the FEL Upgrade installation and commissioning activities with budget estimates for the remainder of FY03 for discussion with the DOE and ONR Program Offices.

A white paper is in preparation for presentation to a subcommittee of the Basic Energy Sciences Advisory Committee on Feb. 22-23 on potential scientific and technical studies with the FEL Upgrade.

WBS 3 (Beam Physics):
Analysis of dipole measurements continues. Both GX and GQ results appear remarkably like Tom Schultheiss' model. Field quality is generally good and we are developing "cuts" of the data that suggest the magnet performance is, or can be made to be, acceptable.

To do:

1. Certify field integrals along design orbit at 80 and 145 MeV/c (initial GX results in and looking good, GQ results underway)

2. Certify field flatness across design orbit at 80 and 145 MeV/ ("Tennant plots" look good for GX, but display some end features in GQ; specific cuts across design orbit are yet to be evaluated)

3. Understand end effects in GQ dipole - ascertain if they are of importance (i.e., complete steps 1 & 2) and if so, perform corrections (see WBS Biallas)

Magnets appear mechanically and electrically reproducible; GQ end corrections will therefore be based on measurements of the next GQ pair, with corrections made in the vault to the pair that has already been measured.

WBS 4 (Injector):
The Gun HVPS and limiting resistors are fully operational. See the note under "Commissioning Activities" above for this week’s high voltage conditioning activities on the gun.

WBS 6 (RF):
All RF equipment is Operational. The final calibration factors will be checked again for Zone 4.

WBS 8 (Instrumentation):
The gun power supply analog board was calibrated from the power supply through to EPICS. Also the HVPS permits (OFF & NOT ON) from the injector area to the DLPC were fully tested and are operational. This allows the drive laser to be operated independent of the Machine Protection System (MPS) when the machine is NOT in beam permit.

The remaining hardware associated with the MPS has been installed and is being commissioned. About 90% of the external inputs have been connected. These include all of the beam viewers, vacuum valves, and some of the interlocks from the dump water skid and raster power supplies. This system checkout will continue for the next few weeks then a formal certification document will be created and the system firm ware will be frozen.

The following work has been done on the video system: The injector cathode camera and brewster window camera have been fully tested and implemented into the system and are operational. The 2nd switcher has been brought online bringing the operational channel count to 128. The 1st switcher chassis had an internal wiring failure that affected the 3rd bank of inputs(channels 33 to 48). This is fixed. Channel 50 has been determined to be dead and is labeled in software until it is fixed. The online video pages have been updated as follows: 1) the upgrade web-cam has been turned into a "recent snap-shot only" page. 2) A non-switchable video input has been set-up to provide the remote-monitoring of commissioning activities (i.e. gun vacuum and optics vacuum gauges). 3) switchable video channels 15 & 16 have been made available online. These signals may all be viewed off-site at: http://laser.jlab.org/video.

The HVPS permits (OFF & NOT ON) from the injector area are fully tested and Operational. The time and date have been corrected on the scope video quad. All of the Superscanner VGA-NTSC converters have been adjusted for position and contrast, including the drive laser autocorrelator video signal. The spiricon video input was restored as 'Sally PC video'. VCR#1 and VCR#2 in the FEL control room are both integrated into the video system again. A 3rd video mux chassis was installed. Once the chassis is fully in-system tested, the video system will be 192 channels. Since the video system has received so much attention recently, this weeks pictures are various views of the heart of the system, the 64:16 video cross-point chassis. These are the chassis that process all the video and AMS signals for the FEL.

An X-term has been permanently installed in the tunnel. It is on a cart in the injector area. This will aid in commissioning components by allowing local operation.

All magnets in the LINAC have been connected to power. Testing is in progress and should be completed by the end of the week. BPMs and magnets in the 3F region are being hooked up and checked as well. Software testing of the Apogee Bipolar power supply, VME controller and interface module has begun with good results. Voltage/current set points and read backs all operated accordingly. The next step is to test multiple VME PS controllers on one IOC for possible interrupt conflicts. Work is near completion for the 1st Arc String shunt modules. Commissioning of the Injector and 1st Linac trim racks has begun.

Vacuum system status: Injector area - all cables have been terminated and connected. All ion pumps have been powered but power to the gate valves has not been connected yet as work in the region still continues. Faults in the Injector area are the position indicator of gate valve VBV0F02A has to be remounted. Ion pump VIP0F02 needs to be replaced as it is contaminated. 1F linac area - All cables have been terminated and connected. All ion pumps have been powered but power to the gate valves has not been connected yet. 2F linac area - All cables have been terminated and connected. All ion pumps have been powered but power to the gate valves has not been connected yet. Exception to this area is the arc region which have not been terminated or routed yet. 3F linac area - All cables have been terminated and connected. All ion pumps have been powered but power to the gate valves has not been connected yet. 4F linac area - All cables have been terminated up to VIP4F09 except for the optical chicane which is still under construction. None of the 4F region has been powered up yet as it is still in construction and not under vacuum. 5F linac area - All cables have been terminated and connected. All ion pumps have been powered but power to the gate valves has not been connected yet, exception to this is the arc region which has not been terminated or routed yet. 1V Optical area - Cables run but not terminated as of yet. Waiting on construction of that area and to be pulled under a vacuum. Vacuum Racks; FL05B21 rack has been completely terminated and will be going through a system checkout in the beginning of next week. FL05B20 rack has had the MPS chassis terminated and verified, also most of the IPPS terminated and operational. FL05B19 rack is now being terminated and is about 50 % complete. Gate Valves and Convectron gauges still to be terminated. System check out of this rack will try to be started toward the end of next week for the components installed. FL05B18 rack is complete as far as hardware, but requires check out of epics signals from the vacuum crates as well as getting the Digitels reading correctly in EPICS. Installation of the 2G00 and 2G00A beamviewer assemblies has begun. Low level checkout of the beamviewer chassis MPS interlock was performed successfully.

System prints, chassis and wiring prints check prints for the HVPS interlock system have been received from EECAD and are under review. Check prints for the following drawings have been received from EECAD and are awaiting review: 05514-D-0004 (Rev.A)-Dump Water Conductivity Monitoring Chassis Assembly Details; 05514-D-0003 (rev.A)-Dump Water Conductivity Monitoring Chassis Wiring Diagram; F0180-F0184-HVPS Interlock Chassis Fabrication and Assembly; F0189-Dump Current Buffer Driver Board Schematic (Version 2); F0185-F0188-Dump Current Buffer Driver (Version 1) Schematic, Wiring Diagram, Fabrication and Assembly Details.

The new ODH Assessment has been completed and is in routing for signature. This will return the vault to ODH 0 status during normal operations as long as OSP #A-00-002-OSP which locks open the return valves for the helium reliefs remains in effect. The upper level and control room can be restored, i.e. the door and ceiling tiles replaced and the ventilation louvers shut, once the penetrations in the floor have been stuffed with fire blocks.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
• No further progress
Optical Chicane Dipoles (GW)
• No progress.
Injector Dipoles (DU/DV)
• No action
Arc 180 Degree Dipoles (GY)
• Don Bullard continued putting together the semicircular-tracked magnet measurement apparatus. He also is working on a few details of the magnet that require TLC before we can measure. The field clamp over the coil crossover wings on the left side interferes so the field clamp has to be extended. The shop is working on this. Some of the GG coil electrical leads were not brazed in the right position and had to be re brazed. On those same leads, the conductor was cut off short on some. Thus the water tubes coming out of the coil conductor come up short of the bulkhead fitting. Don will drill out the bulkhead to allow the compression fittings to work. The design position of the LCW manifolds for the hoses to the GG coils doesn’t allow good hose placement. Don is fabricating extensions to the manifold bracket that allow a new position for the manifolds with a better hose placement.
• Wang is potting the first coil of the second GY next. They predict shipment in three weeks.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
• David Douglas allowed the installation of the second Reverse Bend (GQ) and the first Bend (GX) into the First Arc after successful measurements.
• The second GX developed a one drop per thirty seconds leak in its bottom coil that is coming out of the bulk of the conductor at the lead. Since this coil is wound from single lengths of hollow copper conductor with no internal joints, (confirmed by a call to Wang NMR) I have to conclude that the leak is coming from a defect in the copper which we supplied. Fortunately Wang NMR has two of the four lengths of copper required and can harvest the remaining two lengths from the 16 over-long lengths within the first GY coil (that was considered not able to be fixed). Wang NMR will charge us the regular price of a GX coils plus the cost of the time to unwind the GY coil copper and re insulate it. We are ordering another coil.
• Wang NMR is predicting shipment of both complete GQ magnets today or Monday
• Analysis of the measurements of the GQs and GX by the integral over the nominal beam arc is continuing.
• We will measure the remaining three GX magnets over the first three days of next week to qualify them for insertion in the machine. We will insert the missing one into the first arc. The remainder will be able to go into the second arc as the installer time is available and after the leaky GX (going in the first dipole position in the second arc) has its coil changed out with the replacement from Wang NMR.
Trim Quad (QT)
• The trim coils to create the GC vertical corrector function within the four trim quads were too small and new ones were ordered. We will get them from Milhous Control on Monday and they will be inserted (making the magnets available for installation) by the end of next week.
Sextupole (SF)
• No progress
Octupole (OT)
• The procurement process for the cores continued with the bids due today.
Beam Line and Vacuum
• Our shop continued working on the last chamber in Optical Chicane. They forecast completion on the 4thth of February.
• The magnets of the Optical Chicane were wired up to the Arc 1 power supply.
• The two 180* Bend chambers continue to await the receipt of the diagnostic windows from Ceramaseal. The latest projected delivery date is late February.
• The Vacuum Group connected girders in the regions around the optical klystron in their drive to button up the machine.
• Most of the correctors and skew quads in the Linac were tested for powering and polarity. Several powering problems are being hunted down.

WBS 11 (Optics):
This week work on the UHV-compatible wiring was completed. When connected to the controls systems, all mirror motion control and their readbacks checked out. A big thanks to the I&C group for their assistance. Work now concentrates on the outcoupler. In the HR vessel, we reconfigured various ports to be more compatible with remote commissioning: a more permanent arrangement for hot nitrogen purging, viewports to view cavity optics and utilities motion, and the pumpout/letup port. We will install the 220 L/sec ion pump today or Monday. The initial "dirty" assembly of the fabricated parts for the diagnostic viewer/laser safety shutter was completed, the parts were UHV cleaned and are ready for installation. Initial assembly of the insertable mirrors after the outcoupler is looking good so far; we learned early this week that one important component isn't released because accelerator hardware takes precedence. This cuts into the float in the schedule, but we have a work-around if it continues to be delayed.

Other commissioning activities:
Drive laser optical transport system alignment was performed last Saturday. This was an excellent end-to-end checkout of that particular subsystem. Final optimization will occur once the gun valve can be opened and laser light can illuminate the photocathode. Work on mounting and aligning the fiber optics for the THz diagnostic was discontinued until we can get fiber with an angled input; the flat end feeds back into the ultrafast laser and terminates mode-locking.

A set up using a grating pair arranged in a 4-pass configuration provided the same amount of pulse-broadening (from 85 fs to 5 ps), but with none of the feedback problems. It is also less sensitive to input beam alignment. Work will begin today on aligning and testing the optical diagnostics we will field in the FEL vault to measure the laser properties.